# The Geometric View of Theories

### v1

@@ 3.2 The semantic conception of theories, dualities and scientific realism
The semantic conception of theories, as originally formulated, does not envisage structures like the ones we have here discussed for the model bundle. For that reason, there is no answer within the semantic view to the question I asked in the Introduction: Where in physics does the geometric structure come from? Also the recent literature on theoretical equivalence, which sometimes does postulate some structures on the set of models, does not offer an answer to this question (recall Section 1). Likewise, there are no answers to be found in the literature about e.g. the realist interpretation of moduli spaces and the metrics on them. Therefore, to present a theory as a model bundle, a generalization of the semantic conception is required, and many of the basic philosophical questions, such as those regarding theoretical equivalence and scientific realism, are again “on the table”. Formally, the semantic view is then recovered as the trivial model bundle, which is globally the product of the fibre and the base. Furthermore, the semantic view usually does not envisage quasi-dualities, so that the structure group is the identity. Thus the semantic view is the special case of a trivial model bundle, which has no particularly interesting structure. In this Section, I will make some comments about both these formal and interpretative issues. A more detailed discussion is left for future work.
@@
Thus we return to our main question Where in physics does the geometric structure come from? In the model bundle, the key structure comes from the way the models are tied together in fibres above overlapping regions, and this information is provided by quasi-dualities. In the Seiberg-Witten theory, quasi-dualities allow us to focus on the states that can be probed at low energies above a given point in the moduli space. Thus the validity of the low-energy description in a given region of the moduli space determines the branch of the fibre that we are in. In the rest of this Section, I will mostly use the example of the Seiberg-Witten theory to answer the above question.
@@
These effects, as well as the non-triviality of the metric, come from quantum effects on the theory’s quantities, especially in terms of the theory’s Lagrangian. In the Seiberg-Witten theory, the non-triviality of the metric is the result of the infinite series of renormalization effects of the prepotential that is obtained from the low-energy Lagrangian. [^37] In classical gauge theories, moduli spaces are also equipped with metrics. For the original derivation of the moduli space metric, and how it can be obtained from the classical action, see Manton (1982: p. 55). For more discussion, see Atiyah and Hitchin (1988: pp. 12-13). A beautiful derivation of the metric and its Kahler structure for vortices is in Manton and Sutcliffe (2004: pp. 205–212). In quantum cosmology (see Appendix B), the metric comes from the evaluation of the wave function, which is expressible in terms of CFT partition functions that play the role of wave-functions equipped with a metric.
@@
As I have suggested before, the model bundle, as a formal structure, is not a substitute for substantive interpretative discussions of theoretical equivalence and dualities. Nothing in the model bundle guarantees that duals or quasi-duals are automatically theoretically equivalent models, and in general they are inequivalent, because the models’ projections to the base are only partly overlapping. Thus the same semantic and epistemic requirements for theoretical equivalence that have been discussed in the recent literature, are also requirements for the theoretical equivalence of the fibres. [^38] Here, I in particular have in mind the requirement of unextendability (De Haro, 2016), and the interpretationalism-motivationalism contrast by Read and Møller-Nielsen (2020). For a discussion of necessary and sufficient conditions for theoretical equivalence of physical theories, see De Haro and Butterfield (2025: Chapter 12). See also Rickles (2017:pp. 63–66) and Huggett and Wüthrich (2025: Chapter 8). For this reason, the interpretative issues that are interesting here do not mainly concern fibre bundle realism: [^39] For example, the present discussion is separate from the question whether physical degrees of freedom are “really” in a principal bundle or only in a vector bundle. For more on that discussion, see Jacobs (2023: pp. 39-40) and Gomes (2024: pp. 3-4). for presenting a theory as a model bundle is not advancing a naïve or ‘direct’ model bundle realist view. [^40] See North (2021: pp. 4, 218), whose direct realism is distinguished from a naïve reading of the physics directly from the mathematics. Rather, the realism must be cautious, in that substantive logico-semantic analysis is required before ontological conclusions can be drawn from a particular theory formulation. Thus the catious realist takes inter-theoretic relations seriously in addressing ontological questions, both within the model bundle itself (intra-theoretical equivalence), as between different model bundles (inter-theoretical equivalence). [^41] For a more detailed discussion of this cautious realist attitude, see De Haro and Butterfield (2025: Section 13.2). Other discussions of realism in the context of dualities include Rickles (2013: p. 319) and Dawid (2013). Therefore, even though I defend no direct realism about model bundles, lifting the notions of ‘model’ and ‘theory’ one level up and putting models together into a single object like a model bundle is a substantive issue for the cautious realist. It is not merely a structural way of comparing models, but it has ontological consequences.
@@
In the UV completion of the Seiberg-Witten theory in string theory, the moduli space metric has a more direct realist interpretation: not just as the effective metric “seen” by the fields, but as part of a higher-dimensional ambient spacetime metric. This involves a full geometrization of the theory, where the role of the moduli space metric is akin to the role of the ambient spacetime in string theory. The moduli space and its metric are realized as part of the spacetime on which D-branes move and intersect. Without entering into details: the ${\cal N}=2$ supersymmetric Yang-Mills theory whose low-energy limit is the Seiberg-Witten theory gets realized, at higher energies, as the world-volume theory on a set of fivebranes (with five dimensions of space and one of time) intersecting with fourbranes (with four dimensions of space and one of time) along three common spatial directions and time. This common $(3+1)$-dimensional spacetime is the spacetime of the ${\cal N}=2$ supersymmetric Yang-Mills theory. The moduli space of the Seiberg-Witten theory is then realized as two of the additional spatial dimensions of the fivebranes. [^44] For details, see Witten (1997: pp. 454–455), Giveon and Kutasov (1999: pp. 1020-1021), and De Boer and De Haro (2004: pp. 178-179). These papers also discuss the embedding of this D-brane configuration into M-theory, in terms of a single M5-brane. Note that, in the type IIA string realization that I discuss in the main text, the fourbranes are extended in one additional spatial dimension. This is the dimension along which they are suspended between two fivebranes (so that the fivebranes are located at fixed points in this dimension). The part of the spacetime metric on that intersection space is then precisely the Seiberg-Witten metric. [^45] This realist attitude towards the moduli space metric contrasts with e.g. Halvorson (2019: p. 278), who mentions ‘certain “nearness” relations’ between models. His use of scare quotes around “nearness” seems to indicate that distance relations on a space of models are not to be interpreted realistically.
@@ 4 Conclusion
In the finite-dimensional case, there is a manifold or an algebraic variety—a moduli space—with topological and geometric structure as the hallmark of the geometric view. In the infinite-dimensional case, the state space is directly equipped with a metric. In both cases, what were formerly ‘theories’ are models of this larger theoretical structure. The examples, available in the physical literature for philosophical exploration, are numerous. [^46] In addition to the examples given in footnote 34, I here give an incomplete list of examples. A major example is the moduli space of soliton solutions and its relation to dynamics: see Manton and Sutcliffe (2004: pp. 102–108, 202–205). More specifically, the moduli space of monopoles is discussed in Atiyah and Hitchin (1988: pp. 14–20). The moduli space of instantons is discussed in Naber (2011: pp. 370–376). Seiberg (1995) has discussed the moduli space of ${\cal N}=1$ supersymmetric gauge theories, which also enjoy a form of electric-magnetic duality.
@@
In the usual conceptions of scientific theories, there is no clear preference for choosing theories as having either fixed or free parameters: or, in other words, for the level of generality that one chooses one’s theory to have. In this sense, philosophical discussions are usually “blind” to parameters, which have no distinguished role within the theory. By constrast, the geometric view, through its “lifting models one level up” and its being geometrically unified, vindicates theories as being equipped with a moduli space of parameters. De Haro and Butterfield (2025: pp. 29-30) have called this an ‘intermediate’ notion of theory. Parameters are not simply variable real or complex numbers: they are geometrically organised in a space whose geometric structures have physical significance.
The geometric view opens up several additional questions, of which I will here mention three. First, the global information that it encodes is itself a natural candidate for a realist interpretation, and it is related to the emergence of new states. Thus an interesting open question is to characterise emergence in terms of the topological properties of the model bundle. [^47] For a framework for emergence that allows for this kind of treament, see De Haro (2019: pp. 17-19) and De Haro and Butterfield (2025: pp. 506-507). Second, developing a geometric view of theories seems closely tied to understanding issues in quantum field theory and quantum gravity. This is illustrated by the prominence of dualities and quasi-dualities in quantum field theory and quantum gravity research over the past several decades. [^48] For more discussion of this point, see the discussion of successor theories in De Haro and Butterfield, 2025: Section 14.1. Thus the geometric view can help theory development both formally, heuristically and interpretatively. Finally, the geometric view more generally seems closely related to the unifying role of dualities that has been stressed by Dawid (2013) in the context of string theory. Thus the geometric view is closely linked, and bound to cast light on, more specific issues in string theory, such as the string theory landscape (itself an example of the geometric view), compactification, background-independence, and the ontology of string theory.
@@ Acknowledgements
I thank Jeremy Butterfield, Enrico Cinti and James Read for discussions and for their detailed comments on this paper. I also thank Nick Huggett for a discussion of some of these materials. I thank St John’s College, Oxford, for their hospitality and support during the completion of this paper, in the summer of 2025. Finally, I thank the audiences at the places where, since 2023, this work has been presented: Oxford and Milan (2023), Utrecht and Nijmegen (2024), and Gdánsk (2025).
@@ Appendix A The Seiberg-Witten Theory
This Appendix presents the Seiberg-Witten theory as the paper’s main case study of the geometric view. [^49] This is due to space constraints: Appendix B will briefly discuss another example. More examples are given in De Haro and Butterfield (2025) and Cinti and De Haro (2025). In the philosophical literature, the Seiberg-Witten theory has been discussed previously, from the point of view of emergence, in Vergouwen and De Haro (2025): and, from the general point of view of the geometric view, in De Haro and Butterfield (2025). Here, I will focus on a number of aspects that have not appeared in these works, and that motivate my specific proposal in Section 3: namely, the theory’s fibre bundle structure. Section A.1 first introduces the theory, emphasising the monodromies of quantitites. Section A.2 then uses this discussion to motivate the geometric view. Section A.3 discusses the geometric structure on the tangent bundle to the moduli sapce. Section A.4 discusses the models in the fibres.
@@ A.1 A survey of the Seiberg-Witten theory
where $Z=an_{e}+a_{D}n_{m}$ is called the central charge, in the sense that it gives a central extension of the super-Poincaré algebra and contains key dynamical information. $a$ parametrizes the degeneracy of the Higgs field, i.e. the freedom that remains after solving its low-energy equation. This leads to a moduli space of vacua. This degeneracy also appears in the expectation value of the Higgs field, $\langle\phi\rangle$. $a_{D}$ is its magentic dual, and it can be found locally by doing a Legendre transformation of the low-energy Wilsonian effective action of the ${\cal N}=2$ supersymmetric Yang-Mills theory.
@@
The fact that the states $(n_{m},n_{e})$ and the quantities $(a_{D}~a)^{T}$ appear as each others’ mathematical duals, and transform in complementary ways under $\mbox{SL}(2,\mathbb{Z})$, expresses the mathematical duality between states and quantities that we expect more generally. The value $Z$ is obtained by evaluating the quantity $(a_{D},a)^{T}$ on a state with given electric and magnetic charges. This will be important when we set up the bundle structure in the next Section.
At first sight, the action of $\mbox{SL}(2,\mathbb{Z})$ looks like a duality, because it maps a state with charges $(n_{m},n_{e})$ to a state with different charges, $(n_{m}^{\prime},n_{e}^{\prime})$, such that important physical quantities such as the central charge and the mass are preserved. Thus it looks like an isomorphism of state spaces that preserves the quantities. [^52] However, this action of $\mbox{SL}(2,\mathbb{Z})$ falls short of being a duality, for two reasons. First, the states are not simply determined by their charges, but also by their spin quantum numbers: and the spin quantum numbers do not match across the duality, because states with different charges also have different spins, and the theory’s spectrum is not invariant under the exchange. (Since the study of the duality is not my main topic in this paper, in most of this Section, I will set this point aside, and return to it at the end of Section A.4.) For more detailed explanations of why ${\cal N}=2$ is not a duality in this sense, see De Haro and Butterfield (2025: pp. 268-269).
@@ A.2 Motivating the geometric view
The monodromy matrix in Eq. (23) is related to the singularity at $u=\infty$. [^54] For an explanation of why these are not merely coordinate singularities, see De Haro and Butterfield (2025: p. 263). Their existence is related to the impossibility of the moduli space metric to be positive everywhere. Physically, as we will discuss below, the singular behaviour happens when the Wilsonian effective action breaks down, because certain states have not been taken into account. These states can then be studied through the study of the monodromies around the singularities. There are two other singularities in the $u$-plane, [^55] For an explanation of this statement, see Seiberg and Witten (1994a: p. 36) and Bilal (1996: pp. 106-107). at points that are labelled $u=\Lambda$ and $u=-\Lambda$, where $\Lambda$ is a dynamically generated scale of strong coupling, where non-perturbative effects become significant. The associated monodromy matrices are:
@@
As explained in De Haro and Butterfield (2025: pp. 263), this can be interpreted in terms of perturbative and non-perturbative states becoming massless at the singularities. Recall that the Seiberg-Witten theory describes the low-energy regime of supersymmetric Yang-Mills theory. This means that massive states, whose mass is much higher than the dynamically generated scale $\Lambda$, are integrated out and do not contribute to the low-energy spectrum.

### v2

@@ 3.2 The semantic conception of theories, dualities and scientific realism
The semantic conception of theories, as originally formulated, does not envisage structures like the ones we have here discussed for the model bundle. For that reason, there is no answer within the semantic view to the question I asked at the end of the Introduction: Where in physics does the structure on the space of theories originate? Also the recent literature on theoretical equivalence, which sometimes does postulate some structures on the set of models, does not offer an answer to this question (recall Section 1). Likewise, there are no answers to be found in the literature about e.g. the realist interpretation of moduli spaces and the metrics on them. Therefore, to present a theory as a model bundle, a generalization of the semantic conception is required, and many of the basic philosophical questions, such as those regarding theoretical equivalence and scientific realism, are again “on the table”. Formally, the semantic view is then recovered as the trivial model bundle, which is globally the product of the fibre and the base. Furthermore, the semantic view usually does not envisage quasi-dualities, so that the structure group is the identity. Thus the semantic view is the special case of a trivial model bundle, which has no particularly interesting structure. In this Section, I will make some comments about both these formal and interpretative issues. A more detailed discussion is left for future work.
@@
Thus we return to our main question Where in physics does the structure on the space of theories originate? In the model bundle, the key structure comes from the way the models are tied together in fibres above overlapping regions, and this information is provided by quasi-dualities. In the Seiberg-Witten theory, quasi-dualities allow us to focus on the states that can be probed at low energies above a given point in the moduli space. Thus the validity of the low-energy description in a given region of the moduli space determines the branch of the fibre that we are in. In the rest of this Section, I will mostly use the example of the Seiberg-Witten theory to answer the above question.
@@
These effects, as well as the non-triviality of the metric, come from quantum effects on the theory’s quantities, especially in terms of the theory’s Lagrangian. In the Seiberg-Witten theory, the non-triviality of the metric is the result of the infinite series of renormalization effects of the prepotential that is obtained from the low-energy Lagrangian. [^43] In classical gauge theories, moduli spaces are also equipped with metrics. For the original derivation of the moduli space metric, and how it can be obtained from the classical action, see Manton (1982: p. 55). For more discussion, see Atiyah and Hitchin (1988: pp. 12–13). A beautiful derivation of the metric and its Kahler structure for vortices is in Manton and Sutcliffe (2004: pp. 205–212). In quantum cosmology (see Appendix B), the metric comes from the evaluation of the wave function, which is expressible in terms of CFT partition functions that play the role of wave-functions equipped with a metric.
@@
As I have suggested before, the model bundle, as a formal structure, is not a substitute for substantive interpretative discussions of theoretical equivalence and dualities. Nothing in the model bundle guarantees that duals or quasi-duals are automatically theoretically equivalent models, and in general they are inequivalent, because the models’ projections to the base are only partly overlapping. Thus the same semantic and epistemic requirements for theoretical equivalence that have been discussed in the recent literature, are also requirements for the theoretical equivalence of the fibres. [^44] Here, I in particular have in mind the requirement of unextendability (De Haro, 2017), and the interpretationalism-motivationalism contrast by Read and Møller-Nielsen (2020). For a discussion of necessary and sufficient conditions for theoretical equivalence of physical theories, see De Haro and Butterfield (2025: Chapter 12). See also Rickles (2017:pp. 63–66) and Huggett and Wüthrich (2025: Chapter 8). For this reason, the interpretative issues that are interesting here do not mainly concern fibre bundle realism: [^45] For example, the present discussion is separate from the question whether physical degrees of freedom are “really” in a principal bundle or only in a vector bundle. For more on that discussion, see Jacobs (2023: pp. 39–40) and Gomes (2025: pp. 512–514). for presenting a theory as a model bundle is not advancing a naïve or ‘direct’ model bundle realist view. [^46] See North (2021: pp. 4, 218), whose direct realism is distinguished from a naïve reading of the physics directly from the mathematics. Rather, the realism must be cautious, in that substantive logico-semantic analysis is required before ontological conclusions can be drawn from a particular theory formulation. Thus the catious realist takes inter-theoretic relations seriously in addressing ontological questions, both within the model bundle itself (intra-theoretical equivalence), as between different model bundles (inter-theoretical equivalence). [^47] For a more detailed discussion of this cautious realist attitude, see De Haro and Butterfield (2025: Section 13.2). Other discussions of realism in the context of dualities include Rickles (2013: p. 319) and Dawid (2013). Therefore, even though I defend no direct realism about model bundles, lifting the notions of ‘model’ and ‘theory’ one level up and putting models together into a single object like a model bundle is a substantive issue for the cautious realist. It is not merely a structural way of comparing models, but it has ontological consequences.
@@
In the UV completion of the Seiberg-Witten theory in string theory, the moduli space metric has a more direct realist interpretation: not just as the effective metric “seen” by the fields, but as part of a higher-dimensional ambient spacetime metric. This involves a full geometrization of the theory, where the role of the moduli space metric is akin to the role of the ambient spacetime in string theory. The moduli space and its metric are realized as part of the spacetime on which D-branes move and intersect. Without entering into details: the ${\cal N}=2$ supersymmetric Yang-Mills theory whose low-energy limit is the Seiberg-Witten theory gets realized, at higher energies, as the world-volume theory on a set of fivebranes (with five dimensions of space and one of time) intersecting with fourbranes (with four dimensions of space and one of time) along three common spatial directions and time. This common $(3+1)$-dimensional spacetime is the spacetime of the ${\cal N}=2$ supersymmetric Yang-Mills theory. The moduli space of the Seiberg-Witten theory is then realized as two of the additional spatial dimensions of the fivebranes. [^50] For details, see Witten (1997: pp. 454–455), Giveon and Kutasov (1999: pp. 1020–1021), and De Boer and De Haro (2004: pp. 178–179). These papers also discuss the embedding of this D-brane configuration into M-theory, in terms of a single M5-brane. Note that, in the type IIA string realization that I discuss in the main text, the fourbranes are extended in one additional spatial dimension. This is the dimension along which they are suspended between two fivebranes (so that the fivebranes are located at fixed points in this dimension). The part of the spacetime metric on that intersection space is then precisely the Seiberg-Witten metric. [^51] This realist attitude towards the moduli space metric contrasts with e.g. Halvorson (2019: p. 278), who mentions ‘certain “nearness” relations’ between models. His use of scare quotes around “nearness” seems to indicate that distance relations on a space of models are not to be interpreted realistically.
@@ 4 Conclusion
In the finite-dimensional case, there is a manifold or an algebraic variety—a moduli space—with topological and geometric structure as the hallmark of the geometric view. In the infinite-dimensional case, the state space is directly equipped with a metric. In both cases, what were formerly ‘theories’ are models of this larger theoretical structure. The examples, available in the physical literature for philosophical exploration, are numerous. [^52] In addition to the examples given in footnote 40, I here give an incomplete list of examples. A major example is the moduli space of soliton solutions and its relation to dynamics: see Manton and Sutcliffe (2004: pp. 102–108, 202–205). More specifically, the moduli space of monopoles is discussed in Atiyah and Hitchin (1988: pp. 14–20). The moduli space of instantons is discussed in Naber (2011: pp. 370–376). Seiberg (1995) has discussed the moduli space of ${\cal N}=1$ supersymmetric gauge theories, which also enjoy a form of electric-magnetic duality.
@@
In the usual conceptions of scientific theories, there is no clear preference for choosing theories as having either fixed or free parameters: or, in other words, for the level of generality that one chooses one’s theory to have. In this sense, philosophical discussions are usually “blind” to parameters, which have no distinguished role within the theory. By constrast, the geometric view, through its “lifting models one level up” and its being geometrically unified, vindicates theories as being equipped with a moduli space of parameters. De Haro and Butterfield (2025: pp. 29–30) have called this an ‘intermediate’ notion of theory. Parameters are not simply variable real or complex numbers: they are geometrically organised in a space whose geometric structures have physical significance.
The geometric view opens up several additional questions, of which I will here mention three. First, the global information that it encodes is itself a natural candidate for a realist interpretation, and it is related to the emergence of new states. Thus an interesting open question is to characterise emergence in terms of the topological properties of the model bundle. [^53] For a framework for emergence that allows for this kind of treament, see De Haro (2019: pp. 17–19) and De Haro and Butterfield (2025: pp. 506–507). Second, developing a geometric view of theories seems closely tied to understanding issues in quantum field theory and quantum gravity. This is illustrated by the prominence of dualities and quasi-dualities in quantum field theory and quantum gravity research over the past several decades. [^54] For more discussion of this point, see the discussion of successor theories in De Haro and Butterfield (2025: Section 14.1). Thus the geometric view can help theory development both formally, heuristically and interpretatively. Finally, the geometric view more generally seems closely related to the unifying role of dualities that has been stressed by Dawid (2013) in the context of string theory. Thus the geometric view is closely linked, and bound to cast light on, more specific issues in string theory, such as the string theory landscape (itself an example of the geometric view), compactification, background-independence, and the ontology of string theory.
@@ Acknowledgements
I thank Jeremy Butterfield, Enrico Cinti, and James Read for discussions and for their detailed comments on this paper. I also thank Silvester Borsboom, Benjamin Feintzeig, and Nick Huggett for discussions of some of these materials. I thank St John’s College, Oxford, for their hospitality and support during the completion of this paper, in the summer of 2025. Finally, I thank the audiences at the places where, since 2023, this work has been presented: Oxford and Milan (2023), Utrecht and Nijmegen (2024), and Gdánsk and Stockholm (2025).
@@ Appendix A The Seiberg-Witten Theory
This Appendix presents the Seiberg-Witten theory as the paper’s main case study of the geometric view. [^55] This is due to space constraints: Appendix B will briefly discuss another example. More examples are given in De Haro and Butterfield (2025) and Cinti and De Haro (2025). In the philosophical literature, the Seiberg-Witten theory has been discussed previously, from the point of view of emergence, in Vergouwen and De Haro (2025): and, from the general point of view of the geometric view, in De Haro and Butterfield (2025). Here, I will focus on a number of aspects that have not appeared in these works, and that motivate my specific proposal in Section 3: namely, the theory’s fibre bundle structure. Section A.1 first introduces the theory, emphasising the monodromies of quantitites. Section A.2 then uses this discussion to motivate the geometric view. Section A.3 discusses the geometric structure on the tangent bundle to the moduli space. Section A.4 discusses the models in the fibres.
@@ A.1 A survey of the Seiberg-Witten theory
where $Z=an_{e}+a_{D}n_{m}$ is called the central charge, in the sense that it gives a central extension of the super-Poincaré algebra and contains key dynamical information. $a$ parametrizes the degeneracy of the Higgs field, i.e. the freedom that remains after solving its low-energy equation. This leads to a moduli space of vacua. This degeneracy also appears in the expectation value of the Higgs field, $\langle\phi\rangle$. $a_{D}$ is its magnetic dual, and it can be found locally by doing a Legendre transformation of the low-energy Wilsonian effective action of the ${\cal N}=2$ supersymmetric Yang-Mills theory.
@@
The fact that the states $(n_{m},n_{e})$ and the quantities $(a_{D},a)^{T}$ appear as each others’ mathematical duals, and transform in complementary ways under $\mbox{SL}(2,\mathbb{Z})$, expresses the mathematical duality between states and quantities that we expect more generally. The value $Z$ is obtained by evaluating the quantity $(a_{D},a)^{T}$ on a state with given electric and magnetic charges. This will be important when we set up the bundle structure in the next Section.
At first sight, the action of $\mbox{SL}(2,\mathbb{Z})$ looks like a duality, because it maps a state with charges $(n_{m},n_{e})$ to a state with different charges, $(n_{m}^{\prime},n_{e}^{\prime})$, such that important physical quantities such as the central charge and the mass are preserved. Thus it looks like an isomorphism of state spaces that preserves the quantities. [^58] However, this action of $\mbox{SL}(2,\mathbb{Z})$ falls short of being a duality, for two reasons. First, the states are not simply determined by their charges, but also by their spin quantum numbers: and the spin quantum numbers do not match across the duality, because states with different charges also have different spins, and the theory’s spectrum is not invariant under the exchange. (Since the study of the duality is not my main topic in this paper, in most of this Section, I will set this point aside, and return to it at the end of Section A.4.) For more detailed explanations of why ${\cal N}=2$ is not a duality in this sense, see De Haro and Butterfield (2025: pp. 268–269).
@@ A.2 Motivating the geometric view
The monodromy matrix in Eq. (23) is related to the singularity at $u=\infty$. [^60] For an explanation of why these are not merely coordinate singularities, see De Haro and Butterfield (2025: p. 263). Their existence is related to the impossibility of the moduli space metric to be positive everywhere. Physically, as we will discuss below, the singular behaviour happens when the Wilsonian effective action breaks down, because certain states have not been taken into account. These states can then be studied through the study of the monodromies around the singularities. There are two other singularities in the $u$-plane, [^61] For an explanation of this statement, see Seiberg and Witten (1994a: p. 36) and Bilal (1996: pp. 106–107). at points that are labelled $u=\Lambda$ and $u=-\Lambda$, where $\Lambda$ is a dynamically generated scale of strong coupling, where non-perturbative effects become significant. The associated monodromy matrices are:
@@
As explained in De Haro and Butterfield (2025: p. 263), this can be interpreted in terms of perturbative and non-perturbative states becoming massless at the singularities. Recall that the Seiberg-Witten theory describes the low-energy regime of supersymmetric Yang-Mills theory. This means that massive states, whose mass is much higher than the dynamically generated scale $\Lambda$, are integrated out and do not contribute to the low-energy spectrum.